# Hawking radiation from the double copy

### v1

@@ Introduction.
It is natural to ask, for example, how gauge theory can contain, or give rise to, gravitational structures which seemingly have no gauge theory equivalent. Black holes are an obvious target of study – how is information on the horizon, boundary conditions, and Hawking radiation encoded in the gauge theory? It is known that the Schwarzschild metric can be obtained from the Coulomb potential through ‘classical double copy’, which more broadly relates exact solutions of the classical Einstein and Maxwell equations Monteiro et al. (2014); Luna et al. (2015); Berman et al. (2019); Alfonsi et al. (2020); Bahjat-Abbas et al. (2020); Adamo and Kol (2022); Luna et al. (2022); White (2024); Kim (2025); Chawla et al. (2024), and within this scheme procedures exist for identifying horizons Chawla and Keeler (2023). What about the quantum theory, though? To what extent is information on black hole physics accessible through amplitude calculations? We will address this question here.
Combining amplitudes-based approaches Aoude and Ochirov (2023); Aoude et al. (2024); Aoki et al. (2025) with the worldline formalism Affleck et al. (1982); Bern and Kosower (1991, 1992); Strassler (1992); Edwards and Schubert (2019); Kälin and Porto (2020); Mogull et al. (2021), we will show here how Hawking radiation and associated structures emerge from double copy. The logic is simple. First, classical geodesics and conserved quantities in Coulomb and Schwarzschild backgrounds obey double copy relations Gonzo and Shi (2021). Second, scattering amplitudes in the semiclassical regime can be calculated via worldline methods which require only knowledge of classical geodesics Hartle and Hawking (1976); Chitre and Hartle (1977); Srinivasan and Padmanabhan (1999); Torgrimsson (2025); Ilderton and Rajeev (2025). This suggests that Hawking radiation should emerge, at all orders in the coupling, by applying what are essentially classical double copy relations to worldline expressions for gauge theory particle creation amplitudes. Focussing on the case of massless particles, we will show that Hawking radiation in a Vaidya spacetime is indeed the double copy of pair creation from a gauge field which is related to Vaidya by classical notions of double copy. As part of this we will also give an exact calculation of particle creation on the gauge theory side, identifying structures which are natural ‘single copy’ analogues of those in gravity, and in existing calculations of Hawking radiation.
@@
For a massless particle, trajectory $z^{\mu}(\tau)$ and charge $gc^{a}$, the colour factor $c^{a}{\tilde{c}}^{a}$ is conserved in the background (2), so it is enough to consider the Abelian problem Gonzo and Shi (2021). We therefore write $c^{a}\tilde{c}^{a}=-qQ$ from here on, in which $q$ and $Q$ are (positive) ‘atomic numbers’ of the Abelian probe particle  [^1] Our conventions are such that what we call a ‘particle’ is attracted to the source, in both Yang Mills and gravity. and background, respectively. The sign in $c^{a}\tilde{c}^{a}$ means that the resulting interaction is attractive. The relevant action for a massless particle in $\sqrt{\mathrm{Vaidya}}$ is then
@@
where $\rm v$}} is again the initial coordinate of the null geodesic at $\mathcal{I}^{\scriptscriptstyle-}$. Comparing (6) and (7) thus suggests that at the level of the solutions of the equation of motion, we have the simple double copy map $qQ\rightarrow M\mathcal{E}$. Shortly, when we discuss scattering amplitudes, we will find it convenient to express quantities in terms of the outgoing energy $\mathcal{E}^{\prime}:=\mathcal{E}+\Delta\mathcal{E}$. To this end we observe that the double copy map could equivalently be stated as
$qQ\rightarrow M\mathcal{E}\quad\textbf{or}\quad qQ\to M\mathcal{E}^{\prime}\,\,\&\,\,\,{{\rm v}}\to{{\rm v}}-{{\rm v}}_{0}\;,$ (8)
where ${{\rm v}}_{0}:=-4GM$, as the latter rules also yield the correct energy difference in Vaidya. Sending ${{\rm v}\to{\rm v}}-{{\rm v}}_{0}$ is motivated by observing that ${{\rm v}}_{0}$ is the initial vv–coordinate for the last radial null geodesic that escapes the collapse in Vaidya, while the corresponding point in $\sqrt{\mathrm{Vaidya}}$  is ${\rm v}=0$. We will confirm below that it is indeed the latter rules in (8) which are obeyed at the amplitude level.
@@
in which $\mathcal{T}$ is the transition matrix, $q$ is the momentum transfer and $b=({{\rm v}},\mathbf{0})$ is the analogue of an impact parameter. We then define eikonal phase $\mathcal{\chi}$ in the usual manner via (momentarily retaining $\hbar$) $e^{i\frac{\chi({{\rm v}})}{\hbar}}\left(1+\mathcal{O}(\hbar)\right)=1+i\tilde{\mathcal{A}}({{\rm v}})$ Di Vecchia et al. (2024). Since variation of $\chi$ with respect to impact parameter leads to the impulse via $\partial_{{\rm v}}\chi({{\rm v}})=-\Delta\mathcal{E}$, we immediately infer $\chi({{\rm v}})$ for $\sqrt{\mathrm{Vaidya}}$ and Vaidya from (6) and (7) respectively. Evidently, it must be the second of (8) that relates the eikonal phases, as these depend on $\mathcal{E}^{\prime}$ and $\rm v$}}. To see this explicitly, we can consider the tree-level contribution to $\chi({{\rm v}})$. The calculations are analogous in Vaidya and $\sqrt{\mathrm{Vaidya}}$, and are discussed along with extensions in Aoude et al. (2024) and Aoude et al. (2025), so we simply quote the two results
@@
where $r_{0}$ is some scale. These expressions are consistent with the latter replacement in (8), but seemingly only modulo the rule ${{\rm v}\rightarrow{\rm v}}-{{\rm v}}_{0}$; however, (10) only holds to leading order in the coupling, and it has already been shown in Aoude et al. (2024) that by including subdominant corrections, the expected dependence of the Vaidya result on ${\rm v}-{{\rm v}}_{0}$ indeed re-emerges.
@@
We immediately observe that the latter double copy map in (8) yields, up to an irrelevant phase,
@@
which is the correct expression for the leading semiclassical approximation of the out-mode in the Vaidya spacetime Ilderton and Rajeev (2025). (It is also the universal form of the same mode near $v\sim{\rm v}_{0}$ in a general radial collapse Hawking (1975).) Thus the semiclassical wavefunctions in $\sqrt{\mathrm{Vaidya}}$  and Vaidya are related by the same double copy map as the classical geodesics  [^2] We note in passing that the former prescription in (8), enforced at the level of the action by $qQ\rightarrow-M\partial_{v}S_{cl}(v)$ in the equation $\partial_{v}S_{cl}(v)=-\mathcal{E}^{\prime}+\Delta\mathcal{E}$, also gives the correct semiclassical action relevant to the mode function (16). It would be interesting to explore the connection between this procedure and the observation in Adamo et al. (2022b) that certain aspects of double copy in plane-wave backgrounds manifest when momentum conserving delta-functions on the gauge-theory side are represented in position-space form.. We can now consider the double copy of amplitudes and observables by following standard procedures to extract the Bogoliubov coefficients from our wavefunctions. These are computed by (Klein-Gordon) overlaps between in- and out-modes evaluated on $\mathcal{I}^{\scriptscriptstyle-}$ Adamo et al. (2024):
@@
The double copy of this mid-time particle creation amplitude (20) is, again using the second of (8),
@@
which is, as claimed, the Hawking radiation amplitude in Vaidya. The associated particle spectrum is thermal Hawking (1974), even though the gauge theory result (20) is not. Note that, unlike for the wavefunctions, the rule $v\rightarrow v-{{\rm v}}_{0}$ does not play a role in the double copy of the amplitude or spectrum (it only introduces a phase factor $e^{i\mathcal{E}{{\rm v}}_{0}}$ into the Bogoliubov coefficients). This is consistent of course, since using (18) in (19) yields (21) directly.
@@ Conclusions.
We have used a novel combination of worldline and amplitudes-based methods to show that Hawking radiation in a collapse metric is the double copy of pair creation in a background gauge field. The metric and gauge backgrounds are themselves related by classical Kerr-Schild double copy. We first worked semiclassically, but to all orders in $G$, showing that mid-time radiation (massless particle creation) in $\sqrt{\mathrm{Vaidya}}$ double copies to Hawking radiation. The appropriate double copy map is inherited from that obeyed by classical geodesics and conserved quantities in the two spacetimes, and this is responsible for introducing the energy dependence which results in a thermal Hawking spectrum. Our results thus tie together three notions of double copy in black hole physics; that for classical solutions, for geodesics, and for amplitudes. This makes it clear that, ultimately, the symmetries of the background, hence conserved quantities, are responsible for the double copy relations between the gauge and gravity theories.
@@ Acknowledgements.
Figure 2: The exact result for the pair creation amplitude $|\alpha^{-1}\beta|$ calculated from (30), in ratio to its asymptotic limit (LABEL:betaalphainverse), as a function of the ratio of energies $\zeta$, for three values of charge $g^{2}qQ$. The asymptotic limit is reached more quickly for larger values of the charge.

### v2

@@ Introduction.
It is natural to ask, for example, how gauge theory can contain, or give rise to, gravitational structures which seemingly have no gauge theory equivalent. Black holes are an obvious target of study – how is information on the horizon, boundary conditions, and Hawking radiation encoded in the gauge theory? It is known that the Schwarzschild metric can be obtained from the Coulomb potential through ‘classical double copy’, which more broadly relates exact solutions of the classical Einstein and Maxwell equations Monteiro et al. (2014); Luna et al. (2015); Berman et al. (2019); Alfonsi et al. (2020); Bahjat-Abbas et al. (2020); Adamo and Kol (2022); Luna et al. (2022); White (2024); Kim (2025); Chawla et al. (2024); Kent and Zimmerman (2025), and within this scheme procedures exist for identifying horizons Chawla and Keeler (2023). What about the quantum theory, though? To what extent is information on black hole physics accessible through amplitude calculations? We will address this question here.
Combining amplitudes-based approaches Aoude and Ochirov (2023); Aoude et al. (2024); Aoki et al. (2025) with the worldline formalism Affleck et al. (1982); Bern and Kosower (1991, 1992); Strassler (1992); Edwards and Schubert (2019); Kälin and Porto (2020); Mogull et al. (2021), we will show here how Hawking radiation and associated structures emerge from double copy. The logic is simple. First, classical geodesics and conserved quantities in Coulomb and Schwarzschild backgrounds obey double copy relations Gonzo and Shi (2021). Second, scattering amplitudes in the semiclassical regime can be calculated via worldline methods which require only knowledge of classical geodesics Hartle and Hawking (1976); Chitre and Hartle (1977); Srinivasan and Padmanabhan (1999); Semrén and Torgrimsson (2025); Ilderton and Rajeev (2025). This suggests that Hawking radiation should emerge, at all orders in the coupling, by applying what are essentially classical double copy relations to worldline expressions for gauge theory particle creation amplitudes. Focussing on the case of massless particles, we will show that Hawking radiation in a Vaidya spacetime is indeed the double copy of pair creation from a gauge field which is related to Vaidya by classical notions of double copy. As part of this we will also give an exact calculation of particle creation on the gauge theory side, identifying structures which are natural ‘single copy’ analogues of those in gravity, and in existing calculations of Hawking radiation.
@@
For a massless particle, trajectory $z^{\mu}(\lambda)$ and charge $gc^{a}$, the colour factor $c^{a}{\tilde{c}}^{a}$ is conserved in the background (2), so it is enough to consider the Abelian problem Gonzo and Shi (2021). We therefore write $c^{a}\tilde{c}^{a}=-qQ$ from here on, in which $q$ and $Q$ are (positive) ‘atomic numbers’ of the Abelian probe particle  [^1] Our conventions are such that what we call a ‘particle’ is attracted to the source, in both Yang Mills and gravity. and background, respectively. The sign in $c^{a}\tilde{c}^{a}$ means that the resulting interaction is attractive. The relevant action for a massless particle in $\sqrt{\mathrm{Vaidya}}$ is then
@@
where $\rm v$}} is again the initial coordinate of the null geodesic at $\mathcal{I}^{\scriptscriptstyle-}$. Comparing (6) and (7) suggests that at the level of the solutions of the equation of motion, we have the simple double copy map $qQ\rightarrow M\mathcal{E}$. However, when we discuss scattering amplitudes, we will find it convenient to express quantities in terms of the outgoing energy $\mathcal{E}^{\prime}:=\mathcal{E}+\Delta\mathcal{E}$. To this end we observe that the double copy map could equivalently be stated as
$qQ\to M\mathcal{E}^{\prime}\quad\&\quad{{\rm v}}\to{{\rm v}}-{{\rm v}}_{0}\;,$} (8)
where ${{\rm v}}_{0}:=-4GM$; this yields the correct energy difference (7) in Vaidya. Sending ${{\rm v}\to{\rm v}}-{{\rm v}}_{0}$ is motivated by observing that ${{\rm v}}_{0}$ is the initial vv–coordinate for the last radial null geodesic that escapes the collapse in Vaidya, while the corresponding point in $\sqrt{\mathrm{Vaidya}}$  is ${\rm v}=0$. We will confirm below that (8) is indeed the rule obeyed at the amplitude level.
@@
in which $\mathcal{T}$ is the transition matrix, $q$ is the momentum transfer and $b=({{\rm v}},\mathbf{0})$ is the analogue of an impact parameter. We then define eikonal phase $\mathcal{\chi}$ in the usual manner via (momentarily retaining $\hbar$) $e^{i\frac{\chi({{\rm v}})}{\hbar}}\left(1+\mathcal{O}(\hbar)\right)=1+i\tilde{\mathcal{A}}({{\rm v}})$ Di Vecchia et al. (2024). Since variation of $\chi$ with respect to impact parameter leads to the impulse via $\partial_{{\rm v}}\chi({{\rm v}})=-\Delta\mathcal{E}$, we immediately infer $\chi({{\rm v}})$ for $\sqrt{\mathrm{Vaidya}}$ and Vaidya from (6) and (7) respectively. Since the eikonal phases in the two theories depend on $\mathcal{E}^{\prime}$ and $\rm v$}}, it seems sensible that they should be related by (8). To see this explicitly, we can consider the tree-level contribution to $\chi({{\rm v}})$. The calculations are analogous in Vaidya and $\sqrt{\mathrm{Vaidya}}$, and are discussed along with extensions in Aoude et al. (2024) and Aoude et al. (2025), so we simply quote the two results
@@
where $r_{0}$ is some scale. These expressions are consistent with (8), but seemingly only modulo the rule ${{\rm v}\rightarrow{\rm v}}-{{\rm v}}_{0}$; however, (10) only holds to leading order in the coupling, and it has already been shown in Aoude et al. (2024) that by including subdominant corrections, the expected dependence of the Vaidya result on ${\rm v}-{{\rm v}}_{0}$ indeed re-emerges.
@@
We immediately observe that the double copy map (8) yields, up to an irrelevant phase,
@@
which is the correct expression for the leading semiclassical approximation of the out-mode in the Vaidya spacetime Ilderton and Rajeev (2025). (It is also the universal form of the same mode near $v\sim{\rm v}_{0}$ in a general radial collapse Hawking (1975).) Thus the semiclassical wavefunctions in $\sqrt{\mathrm{Vaidya}}$  and Vaidya are related by the same double copy map as the classical geodesics  [^2] We note in passing that the prescription $qQ\rightarrow M\mathcal{E}$ mentioned ahead of (8), enforced at the level of the action by $qQ\rightarrow-M\partial_{v}S_{cl}(v)$ where $\partial_{v}S_{cl}(v)=-\mathcal{E}^{\prime}+\Delta\mathcal{E}$, also gives the correct semiclassical action relevant to the mode function (16). It would be interesting to explore the connection between this procedure and the observation in Adamo et al. (2022b) that certain aspects of double copy in plane-wave backgrounds manifest when momentum conserving delta-functions on the gauge-theory side are represented in position-space form.. We can now consider the double copy of amplitudes and observables by following standard procedures to extract the Bogoliubov coefficients from our wavefunctions. These are computed by (Klein-Gordon) overlaps between in- and out-modes evaluated on $\mathcal{I}^{\scriptscriptstyle-}$ Adamo et al. (2024):
@@
The double copy of this mid-time particle creation amplitude (20) is, again using (8),
@@
which is, as claimed, the Hawking radiation amplitude in Vaidya. The associated particle spectrum is thermal Hawking (1974), even though the gauge theory result (20) is not. Note that, unlike for the wavefunctions, the rule $v\rightarrow v-{{\rm v}}_{0}$ does not play a role in the double copy of the modulus of the pair creation amplitude (20), or the spectrum (it only introduces a phase factor $e^{i\mathcal{E}{{\rm v}}_{0}}$ into the Bogoliubov coefficients). This is consistent of course, since using (18) in (19) yields (21) directly.
@@ Conclusions.
We have used a novel combination of worldline and amplitudes-based methods to show that Hawking radiation in a collapse metric is the double copy of pair creation in a background gauge field. The metric and gauge backgrounds are themselves related by classical Kerr-Schild double copy. We first worked semiclassically, but to all orders in $G$, showing that mid-time radiation (massless particle creation) in $\sqrt{\mathrm{Vaidya}}$ double copies to Hawking radiation. The appropriate double copy map is inherited from that obeyed by classical geodesics and conserved quantities in the two spacetimes, and this is responsible for introducing the energy dependence which results in a thermal Hawking spectrum. Our results thus tie together three notions of double copy in black hole physics; that for classical solutions, for geodesics, and for amplitudes. This makes it clear that, ultimately, the symmetries of the background, hence conserved quantities, are responsible for the double copy relations between the gauge and gravity theories. (See Kent and Zimmerman (2025) for a discussion of classical double copy in terms of Killing vectors.)
@@ Acknowledgements.
Figure 2: The exact result for the modulus of the pair creation amplitude $|\alpha^{-1}\beta|$ calculated from (30), in ratio to its asymptotic limit (20), as a function of the ratio of energies $\zeta$, for three values of charge $g^{2}qQ$. The asymptotic limit is reached more quickly for larger values of the charge.